OpenStudy (anonymous):

f(x) = 2cos(x) What is the domain and range? Can you graph this function as well ?

OpenStudy (anonymous):

I think cosine functions go forever, the domain is (- infinity, infinity)

OpenStudy (anonymous):

range is -2,2

OpenStudy (anonymous):

and the domain?

OpenStudy (anonymous):

domain [0,pie] range[-1,1]

OpenStudy (anonymous):

can you graph this for me?

OpenStudy (anonymous):

hold the phone. you can take the cosine of any number. therefore, as meverett said, the domain is \[(-\infty,\infty)\]

OpenStudy (anonymous):

so the domain is all real numbers?

OpenStudy (anonymous):

and the range?

OpenStudy (anonymous):

yes, all real number \[\mathbb R\]

OpenStudy (anonymous):

and the range?

OpenStudy (anonymous):

\[[-2,2]\]

OpenStudy (anonymous):

what zbay said

OpenStudy (anonymous):

can you graph this for me so i can get a better view?

OpenStudy (anonymous):

range = [_2,2]

OpenStudy (anonymous):

click the link i posted it's graphed

OpenStudy (anonymous):

zbay graphed it using wolfram. link is above

OpenStudy (anonymous):

which one is the correct one? the first or second graph?

OpenStudy (anonymous):

they are both correct just looked at in different windows

OpenStudy (anonymous):

i can save it..do you mind if you can draw it?

OpenStudy (anonymous):

cant*

OpenStudy (anonymous):

Right click the graph you want and you should be able to save it that way.